# DetailFlow: 1D Coarse-to-Fine Autoregressive Image Generation via Next-Detail Prediction

### v1

@@ Abstract
This paper presents DetailFlow, a coarse-to-fine 1D autoregressive (AR) image generation method that models images through a novel next-detail prediction strategy. By learning a resolution-aware token sequence supervised with progressively degraded images, DetailFlow enables the generation process to start from the global structure and incrementally refine details. This coarse-to-fine 1D token sequence aligns well with the autoregressive inference mechanism, providing a more natural and efficient way for the AR model to generate complex visual content. Our compact 1D AR model achieves high-quality image synthesis with significantly fewer tokens than previous approaches, i.e. VAR/VQGAN. We further propose a parallel inference mechanism with self-correction that accelerates generation speed by approximately 8× while reducing accumulation sampling error inherent in teacher-forcing supervision. On the ImageNet 256$\times$256 benchmark, our method achieves 2.96 gFID with 128 tokens, outperforming VAR (3.3 FID) and FlexVAR (3.05 FID), which both require 680 tokens in their AR models. Moreover, due to the significantly reduced token count and parallel inference mechanism, our method runs nearly 2× faster inference speed compared to VAR and FlexVAR. Extensive experimental results demonstrate DetailFlow’s superior generation quality and efficiency compared to existing state-of-the-art methods.
@@ 1 Introduction
  • Improved token efficiency. DetailFlow significantly reduces token requirements, achieving 2.96 gFID using a 326M parameter AR model with only 128 tokens and nearly 2× faster inference speed on the ImageNet 256$\times$256 benchmark, compared to 680 tokens required by recent state-of-the-art methods such as VAR tian2024visual and FlexVAR jiao2025flexvar .
@@ 2.2 Visual generation
Autoregressive models. The definition of an appropriate token sequence is crucial for autoregressive models in image generation tasks. Inspired by NLP, the next-token prediction paradigm for sequentially generating discrete image tokens has been naturally extended to decoder-only Transformer architectures. VQGAN esser2020taming , Parti yu2022scaling and EMU3 wang2024emu3 adopt a raster-scan strategy to arrange 2D image tokens into a 1D ordered token sequence by scanning row by row. While this method facilitates token-level prediction in a manner analogous to text generation, alternative approaches have explored more structured generation schemes. Notably, VAR tian2024visual introduces a next-scale prediction framework, which departs from the conventional sequential token prediction by generating coarser-to-finer representations across multiple scales. This hierarchical strategy enables the model to capture global image structure before refining local details.
@@ A.1.2 AR model Setup
Building on the autoregressive LlamaGen architecture, we incorporate learnable mask tokens inspired by PAR [43] to enable parallel decoding. We also use the 2D positional encoding to enhance the model’s ability to distinguish between different token groups and intra-group tokens. As illustrated in Fig. 6(b), the attention pattern of the AR model is structured such that the first group uses causal attention for next-token prediction, while subsequent groups apply bidirectional attention internally and causal attention across groups. This configuration allows for parallel prediction of later groups while preserving autoregressive dependencies. Training is conducted primarily on 16 A800 GPUs.
@@ A.3 Future Work
LlamaGen[36] 2.19 20.79 GPT-L 3.81 248.3 0.83 0.52 343M 256 256 12.58
O.-MAGVIT2[26] 1.17 22.64 AR-B 3.08 258.3 0.85 0.51 343M 256 256 -
PAR[43] 0.94 - PAR-L-4 3.76 218.9 0.84 0.50 343M 576 147 3.38
VAR[38] 0.9 - VAR-d16 3.30 274.4 0.84 0.51 310M 680 10 0.15
FlexVAR[19] - - VAR-d16 3.05 291.3 0.83 0.52 310M 680 10 0.15

### v2

@@ Abstract
This paper presents DetailFlow, a coarse-to-fine 1D autoregressive (AR) image generation method that models images through a novel next-detail prediction strategy. By learning a resolution-aware token sequence supervised with progressively degraded images, DetailFlow enables the generation process to start from the global structure and incrementally refine details. This coarse-to-fine 1D token sequence aligns well with the autoregressive inference mechanism, providing a more natural and efficient way for the AR model to generate complex visual content. Our compact 1D AR model achieves high-quality image synthesis with significantly fewer tokens than previous approaches, i.e. VAR/VQGAN. We further propose a parallel inference mechanism with self-correction that accelerates generation speed by approximately 8× while reducing accumulation sampling error inherent in teacher-forcing supervision. On the ImageNet $\times256$ benchmark, our method achieves 2.96 gFID with 128 tokens, outperforming VAR (3.3 FID) and FlexVAR (3.05 FID), which both require 680 tokens in their AR models. Moreover, due to the significantly reduced token count and parallel inference mechanism, our method runs nearly 2× faster inference speed compared to VAR and FlexVAR. Extensive experimental results demonstrate DetailFlow’s superior generation quality and efficiency compared to existing state-of-the-art methods.
@@ 1 Introduction
  • Improved token efficiency. DetailFlow significantly reduces token requirements, achieving 2.96 gFID using a 326M parameter AR model with only 128 tokens and nearly 2× faster inference speed on the ImageNet $\times256$ benchmark, compared to 680 tokens required by recent state-of-the-art methods such as VAR tian2024visual and FlexVAR jiao2025flexvar.
@@ 2.2 Visual generation
Autoregressive models. The definition of an appropriate token sequence is crucial for autoregressive models in image generation tasks. Inspired by NLP, the next-token prediction paradigm for sequentially generating discrete image tokens has been naturally extended to decoder-only Transformer architectures. VQGAN esser2020taming, Parti yu2022scaling and EMU3 wang2024emu3 adopt a raster-scan strategy to arrange 2D image tokens into a 1D ordered token sequence by scanning row by row. While this method facilitates token-level prediction in a manner analogous to text generation, alternative approaches have explored more structured generation schemes. Notably, VAR tian2024visual introduces a next-scale prediction framework, which departs from the conventional sequential token prediction by generating coarser-to-finer representations across multiple scales. This hierarchical strategy enables the model to capture global image structure before refining local details. Similarly, CART roheda2024cart proposes a "next-detail" prediction strategy that compositionally generates an image by first predicting a "base" factor, which captures the global structure, and then iteratively adding "detail" factors to refine local features.
@@ A.1.2 AR model Setup
Building on the autoregressive LlamaGen architecture, we incorporate learnable mask tokens inspired by PAR [wang2024parallelized] to enable parallel decoding. We also use the 2D positional encoding to enhance the model’s ability to distinguish between different token groups and intra-group tokens. As illustrated in Fig. 6(b), the attention pattern of the AR model is structured such that the first group uses causal attention for next-token prediction, while subsequent groups apply bidirectional attention internally and causal attention across groups. This configuration allows for parallel prediction of later groups while preserving autoregressive dependencies. Training is conducted primarily on 16 A800 GPUs.
@@ A.3 Future Work
LlamaGen[sun2024autoregressive] 2.19 20.79 GPT-L 3.81 248.3 0.83 0.52 343M 256 256 12.58
O.-MAGVIT2[luo2024open] 1.17 22.64 AR-B 3.08 258.3 0.85 0.51 343M 256 256 -
PAR[wang2024parallelized] 0.94 - PAR-L-4 3.76 218.9 0.84 0.50 343M 576 147 3.38
VAR[tian2024visual] 0.9 - VAR-d16 3.30 274.4 0.84 0.51 310M 680 10 0.15
FlexVAR[jiao2025flexvar] - - VAR-d16 3.05 291.3 0.83 0.52 310M 680 10 0.15